OpenStudy (anonymous):

Solve. 3x2 - 11x = 4 my answer: x = -11 and x = 3

OpenStudy (anonymous):

x=-1/3 & x=4

Parth (parthkohli):

Hmm, how did you get that? @InsanelyChaotic

OpenStudy (anonymous):

the - 11 would mean it'd be negative 11?

OpenStudy (anonymous):

hmmm...let me revise..

OpenStudy (anonymous):

you need to get the x by itself on one side and all the numbers on one side.

OpenStudy (anonymous):

-1/3

Parth (parthkohli):

\[3x^2 - 11x - 4 = 0\]Do you know how to solve quadratics by factoring?

OpenStudy (anonymous):

no

Parth (parthkohli):

Then which way do you know?

OpenStudy (radar):

You can always use the quadratic formula. Although, sometimes factoring is faster and easier. I am sure that is how @Blah_blah_blah got the solution so fast.

OpenStudy (radar):

In this case it does factor very easily.

Parth (parthkohli):

@radar: Blah_blah_blah used Wolfram :p

OpenStudy (radar):

LOL!!! O.K.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

they're on to me! quick! away!!!! xp

OpenStudy (anonymous):

hahahahhahaha

OpenStudy (radar):

@InsanelyChaotic your answer in the question box was incorrect, do you still need any assistance?